### v1

# dsld: A Socially Relevant Tool for Teaching Statistics

Abstract
The growing power of data science can play a crucial role in addressing social discrimination, necessitating nuanced understanding and effective mitigation strategies of potential biases. “Data Science Looks At Discrimination” (dsld) is an R and Python package designed to provide users with a comprehensive toolkit of statistical and graphical methods for assessing possible discrimination related to protected groups, such as race, gender, and age. Our software offers techniques for discrimination analysis by identifying and mitigating confounding variables, along with methods for reducing bias in predictive models.
In educational settings, dsld offers instructors powerful tools to teach important statistical principles through motivating real-world examples of discrimination analysis. The inclusion of an 80-page Quarto book further supports users—from statistics educators to legal professionals—in effectively applying these analytical tools to real-world scenarios.
Keywords: Data science education; Statistical learning tools; Algorithmic fairness; Discrimination analysis; Confounding effects; Quarto notebook
1 Introduction
Statistics—the class students love to hate! It’s hard to think of a course less popular, yet required by more majors, than statistics. Recent studies have found a negative student perception of statistical courses, ranging from undergraduate to graduate level course work (Naidu and Arumugam, 2014; Dani and Al Quraan, 2023). Perhaps relabeling as “data science” will help a bit; however, the subject is badly in need of better motivation. To aid this, a number of remedies have been proposed, ranging from the flipped classroom (Kovacs et al., 2021) to stories involving Disney characters (Peters, 2019). The American Statistical Association also has suggestions (Carver et al., 2016).
Our package, dsld (Data Science Looks at Discrimination), takes a different approach by appealing to students’ awareness of social issues (Bowen et al., 2017). The software provides both analytical and graphical/tabular methods for investigating possible bias related to race, gender, age, and other potential sensitive variables. Specifically, two broad categories are addressed:
  • Detection of discrimination: This section focuses on identifying and compensating for confounding variables. For instance, is there a gender gap in wages after taking into account confounders such as age, occupation, number of weeks worked, etc.?
  • Addressing bias in prediction: This section focuses on the reduction of bias in the context of predictive algorithms. For example, consider a tool to aid in granting loan applications. If an applicant’s race is included as a predictor – either explicitly or through proxy variables – how can one mitigate its effect?
The value of the package is greatly enhanced via the use of a tightly integrated open source textbook, written in Quarto (Allaire et al., 2024). The book is not a user manual for the package, but instead serves as a detailed treatment of the statistical concepts themselves, illustrated with dsld examples.
This is no toy. On the contrary, the package will be quite useful in social science research, internal HR analyses, and in discrimination litigation. Both parametric and nonparametric regression models are available.
We note other R packages focusing on analysis of discrimination and related issues: divseg (Kenny, 2022) is concerned with urban racial segregation; genderstat (Arafin Ayon, 2023) is “…an exhaustive tool developed for the R…programming environment, explicitly devised to expedite quantitative evaluations in the field of gender studies;” segregation (Elbers, 2021) is a tool for the calculation of relationships in two-way contingency tables, including with grouping, with a typical intended use case being analysis of urban racial segregation. Several packages address the issue of fairness in prediction, including fairML (Scutari, 2023a); fairmodels (Wiśniewski and Biecek, 2021); and fairness (Kozodoi and V. Varga, 2021).
The organization of the remainder of this paper is as follows: Section 2 introduces the package and Quarto book. Section 3 covers detection of discrimination, and Section 4 addresses reduction of bias in prediction. Finally, the paper concludes with a discussion and future perspectives in Section 5.
Some notation: We have a response variable $Y$ related to a vector of covariates $X$, and a sensitive variable $S$; the latter may be continuous, binary or categorical. $Y$ can be continuous or binary, with coding 1 and 0 in the latter case. In predicting $Y$ in a new case, the predicted value is denoted by $\widehat{Y}$.
2 The dsld Package and Quarto Book
The dsld package was developed in 2023 by seven undergraduates at the University of California, Davis, working under the direction of Norman Matloff. It currently consists of 24 functions, plus an associated open source textbook. It is available on CRAN, and the latest version can be found at https://github.com/matloff/dsld.
2.1 Relation to qeML package
As noted, the package consists of both analytical and graphical/tabular routines. Many of the former functions are wrappers for routines in other packages, adding discrimination-specific functionality.
In the context of using dsld for educational purposes, simplicity of the interface is of high importance to ensure accessibility across a broad student audience. It attains this in many cases via its use of the qeML package. The latter is designed to provide an exceptionally simple, direct user interface. For instance, to fit a linear model to predict mpg in R’s built-in mtcars dataset, in qeML one makes the simple call:
qeLin(data = mtcars, yName = mpg’)
No preparatory calls, no setup of any kind, just an easy request for a model fit. To do the same, but using random forests, just call:
qeRF(data = mtcars, yName = mpg’)
Use qeKNN(data = mtcars, yName = ‘mpg’) for K-Nearest Neighbors, and similar commands for other methods. Flexibility is retained via the number of default arguments.
2.2 Introducing dsld
As noted, many dsld functions wrap functions from other packages while adding material specific to discrimination analysis. To make this concrete, consider dsldLinear, which wraps qeLin (which in turn wraps R’s lm). Here are the call forms of the two functions:
qeLin(data, yName, noBeta0 = FALSE, holdout = floor(min(1000, 0.1 * nrow(data))))
dsldLinear(data, yName, sName, interactions = FALSE, sComparisonPts = NULL, useSandwich = FALSE)
where yName is the name of the response variable Y.
Just as all qeML argument lists begin with data and yName, the same is true with most dsld functions, with a standard third argument being sName.
Furthermore, we see that dsldLinear has some new arguments not present in qeML, which are central to the discrimination analysis:
  • sName: the name of the sensitive variable $S$.
  • interactions: if true, include interactions terms with $S$.
  • sComparisonPts: an argument related to interactions, explained further below.
  • useSandwich: if true, indicates usage of the sandwich method to deal with heteroscedasity (Boe et al., 2023). Not directly related to discrimination analysis.
If interactions = TRUE is set, interactions between the sensitive variable and the predictors/features in our dataset will be modeled. Actually, this approach involves fitting separate linear models for each level of the sensitive variable, which is statistically equivalent to modeling these interactions directly. The presence or absence of interaction terms plays an important role in the package and in the Quarto book. Is this a substantial difference in mean Y, given the co-variates X, across the levels of the R factor S? If so, is this difference uniform with respect to X?
In the case of interactions there is no single “treatment effect” of the sensitive variable. We can no longer speak of “the” difference in mean wage between men and women, as this difference may vary by, say, age. Accordingly, the user can specify several points at which to compare the effects of the different levels of S; estimated differences in mean conditional Y are given for the user-specified points sComparisonPts.
The package includes numerous graphical/tabular functions, some of them wrappers to existing packages, such as freqparcoord, and others standalone. These are used both for preliminary exploration of one’s data, and also for visual illustrations of the results found analytically.
For example, one might use dsldLogit to estimate parameters in a logistic model predicting probability of passage of the bar examination (binary response), based on LSAT scores provided in on the Law School Admissions dataset. We can then supplement these results with graphics such as Figure 1. Suppose we restrict the analysis to students with low GPAs. At initial glance, there appears to be significant racial differences at lower LSAT values, but these differences begin to diminish at higher scores. Notably, the probability curve for Asians shows an unexpected decline at higher LSAT values, warranting additional investigation. It is important to note that these results are purely exploratory, and further analysis is required to draw more definitive conclusions.
Figure 1: Predicted probabilities of passing the bar exam based on LSAT scores, illustrating racial disparities among various groups. Results are limited to students with GPAs below 2.70.
2.3 The Role of Nonparametric Regression Models
As noted in the last section, realistic analyses of discrimination must in many cases account for interaction effects between the covariates $X$ and the sensitive variable S. Often, one must also consider nonlinear effects.
One may address nonlinearity via low-degree polynomial models, but modern computing power is such that a laptop computer can easily handle full nonparametric regression methods, even in fairly large datasets. In dsld, we feature Random Forests (RFs) as our primary non-parametric approach (Hastie et al., 2009).
Note that the statistical term non-parametric regression models corresponds in the computer science community with machine learning (ML), though there is arguably a difference in interpretation, a running debate ever since statistician Leo Breiman’s famous essay, “Statistical Modeling: The Two Cultures,” was published (Breiman, 2001b).
We chose RFs for several reasons. First, it is one of the most widely recognized non-parametric regression methods among statisticians, developed primarily by statisticians in the mid-90s and 2000s (Breiman, 2001a). (Less well known, but also significant, is the work by computer scientists such as Tin Ho Kam and her coauthors (Ho, 1995).) Second, RFs can be easily explained to non-statisticians, due to their “if this, then that” flow chart nature.
2.4 The Quarto book
The Quarto book, about 100 pages in length, is the product of all eight of the package authors. As mentioned, though it uses dsld examples, it presents the statistical concepts, rather than serving as a user manual for the package. Moreover, it emphasizes understanding of the methods beyond just their definition, in practical senses. Below is an excerpt:
One may have specific confounders in mind for a particular analysis, but it is often unclear as to which to use, or for that matter, why not use them all?…
Technically, almost any variable is a confounder. The impact may quite minuscule, but through a long chain of relations among many variables, there will usually be at least some connection, though again possibly very faint…
…there are several issues to consider not using the full set of variables i.e. every variable other than $Y$ and $S$:
  • It may result in overfitting, resulting in large standard errors.
  • It is unwieldy, difficult to interpret. Many treatments of these issues speak of a desire for a “parsimonious” model.
The math involved in the book is minimized, and the material should be accessible to students who have taken a (noncalculus-based) course in elementary statistics. (There are a few optional “starred sections” covering advanced topics.)
2.5 Datasets
Both through the dsld package itself, and through the packages that it wraps, a number of built-in datasets are available to greatly enhance its usefulness as a teaching tool. Note that these datasets tend to be of an observational nature, an issue covered in the Quarto book.
3 Detecting Discrimination
Discrimination is a significant social issue in the United States and many other countries. There is lots of available data with which one might investigate possible discrimination. But how might such investigations be conducted? There is a wide range of classical parametric methods available for this purpose, and non-parametric regression methods have recently gained popularity, particularly in applications like HR management (Frissen et al., 2022). The dsld package offers both graphical and analytical tools to detect potential biases, which are particularly beneficial for students. These tools can enhance understanding and intuition during classroom discussions by connecting them to broader social contexts.
In this section, we will use the lsa dataset on law school admissions, focusing on race as our sensitive variable of interest. The racial categories included are Asian, Black, White, Hispanic, and Other. This dataset is available through the dsld package for further exploration and analysis.
3.1 Graphical/Tabular Methods
Effective graphs and visualizations can significantly aid in students’ understanding of data. The dsld package offers various graphical methods to investigate discrimination issues, including dsldDensityByS, dsldConditDisparity, dsldConfounders and dsldScatterplot3D.
Consider investigating potential discrimination in the context of college and graduate school admissions. There has been growing criticism of standardized testing, arguing it disproportionately favors students with higher family income and other resources (Foreiter, 2021). Indeed, many studies have suggested unfair discrepancies in test results between Black and White students (Dixon-Roman et al., 2013). As a result, many institutions have removed standardized testing requirements like the SAT and GRE from their application processes. This debate underscores the importance of examining biases and motivates our study of the lsa dataset.
dsldConditDisparity
Suppose we want to analyze the effect of LSAT scores on passing the bar examination to investigate potential racial disparities. Similar to the example shown in Figure 1, our analysis now encompasses all students, including those with high GPAs. We can calculate the differences in the probability of passing the bar for a given LSAT score among various racial groups. The function dsldConditDisparity is particularly well-suited for this purpose, as it visualizes such conditional disparities and may provide insights into how LSAT scores and bar examination outcomes may differ among racial groups.
dsldConditDisparity(data = lsa, yName = bar’, sName = race1’, xName = lsat’, condits = NULL)
Figure 2: Predicted probabilities of passing the bar exam based on LSAT scores, illustrating racial disparities among various groups. Results include all students in the lsa dataset.
Notably, all non-White groups initially exhibit similar outcomes in Figure 2, with higher passing probabilities observed for White students up through mid-range LSAT levels. However, as mentioned earlier, all racial groups display comparable passing probabilities at higher LSAT ranges. This observation suggests the potential relevance of racial interaction terms in formal analyses and serves as a valuable starting point for classroom discussions on disparities in educational outcomes based on race.
Note that the dsldConditDisparity includes an optional argument condits (“conditions”). One could restrict the domain of students to those with lower undergraduate GPAs ($\leq$ 2.70), as shown in Figure 1, to conduct similar analyses.
dsldDensityByS
To further investigate the relationship between race and LSAT scores, we can generate a density plot using the dsldDensityByS function. This method enables users to visualize densities of a response variable $Y$ segmented by a sensitive variable $S$, with bandwidth control. This visualization provides insight into potential differences of the distribution of LSAT scores across different racial groups.
This is complementary to our previous graph, Figure 2, which suggested that White students pass the bar exam at higher rates than other races, even at similar LSAT levels. This difference may be exacerbated if White candidates also tend to do better on the LSAT. Consider the following plot:
dsldDensityByS(data = lsa, cName = lsat’, sName = race1’)
Figure 3: Density plot of LSAT scores segmented by race. The bandwidth parameter functions similarly to the bin width in a histogram and allow adjustment of the plot’s granularity. Default bandwidth is set at 1.
The density plot in Figure 3 illustrates the distribution of LSAT scores across different racial groups, highlighting potential bias. However, that apparent effect may be influenced by other confounding variables. Confounders influence both dependent and independent variables, possibly leading to spurious associations (McNamee, 2003). The function dsldConfounders acts similarly, but it is Plotly-based, and thus features interactive user access.
dsldScatterPlot3D
Analyzing relationships among LSAT scores, undergraduate GPA, and family income can reveal potential confounding interactions.We can visualize these relationships by race using dsldScatterPlot3D, which offers a 3-dimensional perspective on multivariable relationships and highlights potential disparities among racial groups in terms of family income and GPA.
[…]
### use a subset of lsa: fam_inc’,‘ugpa’,‘gender’,‘lsat’,‘race1
lsa1 <- lsa[,c(‘fam_inc’,‘ugpa’,‘gender’,‘lsat’,‘race1’)]
dsldFreqPCoord(data = lsa1, m = 75, sName = race1’)
Figure 4: Visualization of the most frequently occurring patterns across different racial groups using parallel coordinates.
Several notable patterns can be seen. Black students exhibit the least variation, with the most common pattern characterized by female students from slightly below-average income families, low undergraduate grades, and significantly low LSAT scores. Patterns for Whites form a mirror inversion of those of Black students, though with greater gender diversity. Asian students display trends similar to Whites. Hispanic students, however, demonstrate a wide range of income levels and grades, yet consistently low LSAT scores.
The graphical analysis tools provided by dsld are invaluable for educators and students in academic settings, aiding in the visualization of complex statistical relationships. For instance, examining LSAT score distributions among different racial groups can initiate discussions on fairness and bias in standardized testing. Similarly, analyzing how family income influences academic outcomes encourages exploration of socioeconomic disparities. These visual methods foster critical thinking by enabling students to interpret real-world data and understand the implications of statistical findings in societal contexts. The Quarto notebook offers a comprehensive analysis of these results and includes relevant insights from other datasets as well.
dsldCHunting
Users can also utilize methods offered by the dsld package to identify potential confounding variables in their own analyses. For instance, the dsldCHunting function employs random forests to evaluate the importance of predictors in predicting $Y$ (without $S$) and $S$ (without $Y$). These predictors are identified as “important predictors” of $Y$ and $S$ and aid in the detection of potential confounding factors. For each $i$ from 1 to intersectDepth, the function reports the intersection of the top $i$ important predictors of Y and $S$. This helps identify potential confounders, with larger values of $i$ indicating inclusion of progressively weaker predictors.
dsldCHunting(data = lsa, yName = bar’, sName = race1’, intersectDepth = 10)
decile3 lsat cluster ugpa decile1 age fulltime gender fam_inc
impForY 0.121 0.063 0.007 0.007 0.006 0.003 0.002 0.001 -0.001
impForS 0.029 0.033 0.023 0.012 0.025 0.003 0.000 0.002 0.006
Table 1: Importance measures for variables in “impForY” and “impForS”. These values indicate feature importance in prediction of $Y$ (without $S$) and $S$ (without $Y$), higher values indicate stronger predictors.
In the context of the lsa dataset, importance measures provide insights into predictors influencing the probability of passing the bar, which may act as confounders concerning race. Table 1 above shows the importance measures: the first row indicates predictors for the probability of passing the bar from other parameters (excluding race), while the second row shows predictors for race from other variables (excluding bar passage). For instance, variables that predict $Y$ (probability of passing the bar) well include decile3, LSAT, and cluster. In predicting $S$ (race), the top predictors are again LSAT, decile3, and cluster. By focusing on variables correlated with both $Y$ and $S$, we can identify potential confounders. For example, the intersection of the top three predictors for $Y$ and $S$ includes decile3, LSAT, and cluster. The analyst has flexibility in choosing the number of confounding variables, ranging from just the top two to more, depending on the individual needs and objectives. Furthermore, results from dsldConfounders can further add to one’s analysis. In particular, this method provides a density plot similar to dsldDensityByS for continuous predictors. For categorical variables, it displays the frequency of the sensitive variable occurring at each level of that variable. The Quarto book offers additional insights of bias detection using various datasets and examples.
3.2 Analytical
While graphical approaches serve as initial steps in exploratory data analysis, this should be followed up with formal methods. the dsld package also provides analytical methods for formal investigation of potential instances of discrimination. Continuing with the LSAT example, suppose we wish to examine pairwise differences in estimated mean LSAT scores across racial groups using a linear model. In this section, we’ll see how dsldLinear facilitates this.
A key issue is whether to include in our model interactions between our covariates $X$ and the sensitive variable $S$, a major point of the dsld package. In educational settings in particular, it is important that students understand that in some cases, there may not be such thing as “the” difference between one group and another; the difference may vary considerably with $X$. We saw this in Figure 2 above, in which $Y$ was an indicator variable for passage of the bar examination.
Here we fit a linear model with $Y$ set to LSAT score, with an eye toward investigating possible bias in the test. Let’s compare the non-interactive and interactive cases:
  • The no-interaction case uses the following call:
    z1 = dsldLinear(data = lsa, yName = lsat’, sName = race1’,
    interactions = FALSE)
    summary(z1)
  • Here is the with-interaction call:
    newData = lsa[c(1,10,100,1000,10000),]
    z2 = dsldLinear(data = lsa, yName = lsat’, sName = race1’,
    interactions = TRUE, sComparisonPts = newData)
    summary(z2)
The key point is that not only does the second call differ from the first in the value of the interactions argument, but also in that the second call includes the argument sComparisonPts. What is happening here? In addition to calculating the usual estimated regression coefficients and so on, dsldLinear, the function facilitates comparison across levels of $S$, as follows.
As noted earlier, in a no-interactions model, we can speak of, for instance, the Black and White difference in mean LSAT scores, independent of the values of $X$. But if interactions of $X$ and $S$ are assumed, there is no notion of “the” Black vs. White difference. That difference will vary according to the values of $X$, so the function estimates the difference in mean $Y$ at values of $X$ that are of interest to the user, as specified in sComparisonPts.
In general, a linear model with interaction terms involving a categorical variable with $m$ levels amounts to fitting $m$ separate linear models.
Thus, in the with-interactions model in the LSAT example, dsldLinear essentially fits five distinct linear models – one for each race.
Let’s take a look at the output of the two models. First, the no-interactions case:
\$Summary Coefficients
Covariate Estimate StandardError PValue
1 (Intercept) 31.98578856 0.448435264 0.000000e+00
2 age 0.02082458 0.005841758 3.641634e-04
3 decile1 0.12754812 0.020946536 1.134602e-09
4 decile3 0.21495015 0.020918737 0.000000e+00
5 fam_inc 0.30085804 0.035953051 0.000000e+00
6 ugpa -0.27817274 0.080430542 5.430993e-04
7 gendermale 0.51377385 0.060037102 0.000000e+00
8 race1black -4.74826307 0.198088318 0.000000e+00
9 race1hisp -2.00145969 0.203504412 0.000000e+00
10 race1other -0.86803104 0.262528590 9.449471e-04
11 race1white 1.24708760 0.154627086 6.661338e-16
\$Sensitive Factor Level Comparisons
Factors Compared Estimates Standard Errors P-Value
1 asian - black 4.748263 0.1980883 0.000000e+00
2 asian - hisp 2.001460 0.2035044 0.000000e+00
3 asian - other 0.868031 0.2625286 9.449471e-04
4 asian - white -1.247088 0.1546271 6.661338e-16
5 black - hisp -2.746803 0.1863750 0.000000e+00
6 black - other -3.880232 0.2515488 0.000000e+00
7 black - white -5.995351 0.1409991 0.000000e+00
8 hisp - other -1.133429 0.2562971 9.764506e-06
9 hisp - white -3.248547 0.1457509 0.000000e+00
10 other - white -2.115119 0.2194472 0.000000e+00
We see for example that the estimated difference in dummy variables for Black and White scores is 6.0000 points with a standard error of 0.1410. In view of the fact that LSAT scores in this dataset ranged from 11 to 48 points, a 6-point difference seems to be very concerning.
On the other hand, the effect of family income is rather small, and indeed gives instructors a chance for a “teachable moment,” as follows. A key concept often emphasized by instructors of statistics courses is the difference between statistical significance versus practical significance. The results here form a perfect example. Consider the variable for family income, which is measured in quintiles, with values 1,2,3,4,5. The coefficient for family income is “very highly significant,” with a p-value of 0 to six figures, yet the effect size is small, about 0.3. The difference between, say, the 3rd and 4th quintiles in terms of mean LSAT is about 0.3 point. That’s minuscule in relation to the general LSAT scores — as mentioned, ranging from 11 to 48 points here — and the Black-White difference of 6 points.
3.3 Causal Models
The function dsldMatchedATE, wrapping Matching::Match offers matched-pairs analysis (Huber, 2023). Direct pairing can be done, or one may opt to use propensity scores, using either glm or qeML::KNN for a nonparametric k-Nearest Neighbors approach. Let’s estimate the “treatment effect” of being female in a possible gender wage gap. We use the svcensus data, which includes information on individuals’ age, education, occupation, income, weeks worked.
data(svcensus)
summary(dsldMatchedATE(svcensus,’wageinc’,’gender’,’male’))
# Estimate 9634.5
# SE……… 380.03
# T-stat….. 25.352
# p.val…… < 2.22e-16
# Original number of observations………….. 20090
# Original number of treated obs…………… 15182
# Matched number of observations…………… 20090
# Matched number of observations (unweighted). 20090
Men are estimated to earn \$9634.50 more than similar women, with a standard error of \$380.03.
Using either logit or k-NN to predict gender, the estimated wage gap becomes:
summary(dsldMatchedATE(svcensus,’wageinc’,’gender’,’male’,
propensFtn=’glm’))
# Estimate 10332
# SE……… 408.39
summary(dsldMatchedATE(svcensus,’wageinc’,’gender’,’male’,
propensFtn=’knn’,k=50))
# Estimate 9877.8
# SE……… 439.73
The function dsldIamb deals with causal discovery. Typically a causal Directed Acyclic Graph (DAG) is merely a reflection of the analyst’s “gut feeling” about relations among the variables. The DAG cannot be derived from the data itself without making some very stringent assumptions (Shalizi, 2024; Scutari, 2023a).
The bnlearn package includes various models of this sort, and dsldIamb wraps the iamb function, with which a DAG may be constructed from the data.
data(svcensus)
# iamb does not accept integer data
svcensus\$wkswrkd <- as.numeric(svcensus\$wkswrkd)
svcensus\$wageinc <- as.numeric(svcensus\$wageinc)
iambOut <- dsldIamb(svcensus)
plot(iambOut)
Figure 5: DAG generated under iamb assumptions
Figure 5 shows the result. Interestingly, the only variable that is not modeled as causal for wage income is gender, though gender is shown as having a weak, non-directional relation with occupation and gender.
Again, the graph is computed under very restrictive assumptions, and the function should be regarded as exploratory.
4 Addressing Bias in Predictive Algorithms
In predicting $Y$ from $X$, one may wish to avoid or limit the influence of $S$. For example, in evaluating an application for a mortgage, we hope that our assessment does not discriminate against women or minorities. Such issues have been the subject of intensive research in recent years. There is even a conference devoted to this effort, the ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT). Here, we present some of dsld’s capabilities in this regard.
The “Hello World” example for this area of research involves the COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) algorithm developed by Northpointe, a commercial entity. The algorithm was designed to predict a defendant’s likelihood of recidivism, thus aiding judges in determining sentences. Note that, as a commercial product, the algorithm is a “black-box,” of completely unknown details. The algorithm has faced criticism following an investigation by the publication ProPublica, which suggested that the algorithm exhibited bias against Black defendants (Angwin et al., 2016). Northpointe has contested these findings, arguing that ProPublica’s analysis is flawed. ProPublica has issued a rejoinder, but in any case, the debate over COMPAS underscores the critical need to address fairness.
Beyond the legal system, predictive methods have rapidly expanded into various commercial applications in recent years, including cybersecurity, healthcare, and e-commerce (Sarker, 2021). These models play a pivotal role in decision-making processes with significant consumer impact. This rise has highlighted concerns about algorithmic fairness in predictive methods (Wehner and Köchling, 2020; Chen, 2023).
Broadly speaking, “fairness” in this context refers to minimizing the influence of sensitive attributes (e.g., race, gender, religion, etc.) on an algorithm’s predictions (Oneto and Chiappa, 2020).
There are several main issues:
  • Defining unfairness: Several measures have been proposed.
  • Reducing unfairness: For a given algorithm, how can we ameliorate its unfairness while maintaining an acceptable predictive accuracy level?
  • Dealing with proxies: Typically there will be proxy variables which, through correlation with S, can result in the latter influencing $\widehat{Y}$ even if S is omitted from the analysis entirely.
The dsld package provides methods for fair predictions with an intuitive, user-friendly interface for various fairness-constrained modeling approaches. In this section, we will use the COMPAS and lsa datasets to the showcase tradeoff between fairness and accuracy on both regression and classification settings. For the COMPAS dataset, we will predict the probability of recidivism with race as the sensitive feature. For the lsa dataset, we will predict LSAT scores with race as the sensitive variable. Both datasets are accessible via the dsld package for further exploration and analysis.
4.1 Relevant Methods Provided by dsld
The dsld package provides wrappers for several functions from the faiML package (Scutari, 2023a) and for the Explicitly Deweighted Features (EDF) methods developed in Matloff and Zhang (2022). Let’s consider the fairML functions first.
FairML
Komiyama et al. (2018), Zafar et al. (2017) and Scutari (2023a) apply ridge-regression-like approaches to reducing the influence of $S$ on the predicted value, $\widehat{Y}$. All these methods are implemented in the fairML package (Scutari, 2023b), with dsld providing interfaces.
Marco Scutari’s work in the fairML package offers a statistical approach to fair machine learning, building on these previous efforts by imposing a ridge-regression-like penalty on the coefficient(s) of S. For example, let’s fit a Fair Generalized Ridge Regression model — a fair adaptation of the Generalized Linear Models, and make a prediction of recidivism in the COMPAS dataset. As an example case, take someone like observation 188 in our dataset: age 23, no prior convictions, male and Black.
cps <- compas1[,c(‘age’,‘priors_count’,‘sex’,‘two_year_recid’,‘race’)]
z <- dsldFgrrm(data=cps, yName=‘two_year_recid’, sName=‘race’, unfairness=0.1)
newx <- cps[188,-4]
predict(z,newx)
# 0.508936
The unfairness argument must be a value in (0,1], with values near 0 being nearly perfectly fair. This controls the Fairness-Utility Tradeoff. Here, the algorithm reports an estimated probability of recidivism of about 51%. What if we had placed no restriction on the role of race?
z <- dsldFgrrm(data=cps, yName=‘two_year_recid’, sName=‘race’,
unfairness=1.0)
predict(z,newx)
# 0.5723716
The probability of recidivism now jumps to 57%.
EDF-Fair
The dsld package also includes wrappers for functions from the work on Explicitly Deweighted Features (Matloff and Zhang, 2022). This approach omits $S$ entirely and mitigates the impact of potential proxy variable through user-specified hyperparameters. The latter enables the user to select a desired point on the Fairness-Utility Tradeoff spectrum. The package offers implementations of linear/logistic regression (works like ridge regression, except with different “$\lambda$” values for each proxy variable), random forests (different proxy variables have different probabilities of being used for splitting a node), and k-nearest neighbors (k-NN) (Euclidean distance but with different weights for each proxy variable).
Let’s apply k-NN on the same example as above.
z <- dsldQeFairKNN(data=cps, yName=’two_year_recid’,
sName=‘race’, deweightPars= list(age=0.5,priors_count=0.1),
yesYVal=‘Yes’)
predict(z,newx)
# 0.52
Here, we entirely omit the race variable from our dataset and deweight the proxies ‘age’ and ‘priors count’ by 0.5 and 0.2 (a weight of 1.0 means no deweighting) to reduce their corresponding predictive power.
4.2 Assessing Fairness and Utility
One possible measure of fairness is to require that $\widehat{Y}$ and S be statistically independent (Johndrow and Lum, 2019). However, this eliminates any use of proxies and thus the ability to achieve a Fairness-Utility Tradeoff.
A broader view would be to ask that $\widehat{Y}$ and S have low correlation. See for instance Li et al. (2023), Kozodoi et al. (2022), Deho et al. (2022), Mary et al. (2019), and Baharlouei et al. (2019) apply work from Lee et al. (2022) and Roh et al. (2023).
The dsld package continues this theme:
  • Measuring Accuracy: To assess utility, we use a holdout set to compute test accuracy from predictions generated by our algorithms. In regression, accuracy is defined as mean absolute prediction error (MAPE); for classification, we use misclassification rate.
  • Measuring Fairness: The measure of fairness is the correlation between $\widehat{Y}$ and $S$, computed using the Kendall Tau statistic. This choice is preferred over the more commonly used Pearson correlation due to its flexibility with both continuous and categorical inputs. While Pearson is geared towards continuous numeric variables, Kendall Tau is also applicable for binary or ordinal integer-valued variables.
4.3 Proxy Hunting
A proxy is a variable that indirectly infers a protected attribute, potentially introducing bias in decision-making even when the protected attribute omitted from the dataset. For example, race or religion may be linked to a city or neighborhood in a city (Morse et al., 2020).
To detect potential proxies in a dataset, the dsld package provides methods, such as dsldOHunting, to help users make informed decisions. Specifically, this function calculates the Kendall Tau correlation between the sensitive feature and all other features in the dataset. Consider use of dsldOHunting on the COMPAS example presented earlier:
dsldOHunting(data = cps, yName = two_year_recid’, sName = race’)
age priors_count sex.Female sex.Male
# race.African-American -0.156 0.175 -0.0414 0.041
# race.Asian 0.016 -0.030 -0.0216 0.021
# race.Caucasian 0.147 -0.100 0.0686 -0.068
# race.Hispanic 0.022 -0.068 -0.0261 0.026
# race.Native American 0.002 0.023 -0.0065 0.006
# race.Other 0.002 -0.087 -0.0129 0.012
The results suggest, say, that “age” and “prior convictions count” might be treated as proxies for being African-American, and possibly de-weighting them accordingly if one uses EDF methods.
The diverse functions offered by dsld provide a range of fair modeling techniques, allowing students to thoroughly explore the balance between fairness and accuracy in prediction tasks. Students are encouraged to apply principles of fair machine learning and illustrate the fairness-utility trade-off using different datasets.
5 Discussion
In this paper, we introduce “Data Science Looks at Discrimination” (dsld) as a powerful tool for educating students in statistics through the application of discrimination analysis and fair machine learning. The software includes a range of analytical and graphical tools that enable students to explore and visualize potential sources of bias and discrimination. The fair machine learning wrappers offer several fairness-constrained machine learning methods, facilitating the quick and seamless deployment of bias mitigation algorithms. Additionally, the accompanying Quarto notebook provides a comprehensive foundation in relevant statistical principles, featuring real-world examples that apply methods provided from dsld, and requires only a basic understanding of mathematics.
As machine learning becomes an increasingly important tool for statistical analysis and prediction, it is essential for users to understand the risks associated with its misapplication. The dsld package not only bridges the gap between traditional and fairness-constrained machine learning algorithms but also serves as an effective educational resource. It equips both current and future students with the knowledge and tools necessary to responsibly apply practical machine learning systems that balance fairness and utility.
SUPPLEMENTARY MATERIAL
The package can be accessed here: https://github.com/matloff/dsld/tree/master. All relevant datasets, functions, and implementation details can be found through the Github repository.